OpenStudy (anonymous):

Solve for m: w = m - 5L. A. m = w - 5L B. m = w + 5L C. m = (w - 5)/L D. m = w/L + 5

OpenStudy (hoblos):

w = m - 5L add 5L for both sides w+5L = m so...m = w + 5L